All Notes (7)
Fire Back on again for remainder of mission.
+190 cc climb, save a little power.
Turning fire off, battery dropping rapidly. Will turn on a segment tonight.
Climbs were set to 20 deg from last mission, changed to 25, should help a bit with currents.
Deployed by the Palmer crew.
Sampling FIRE 1 yo/segment. Heading to UD cross calibration location and then will run Filipa's zig zag from NE Palmer canyon to the SW
Optode removed for 2015-01-05 LTER deployment
Equipped with SB:
CTD SN 0139
FIRE SN 01
OPTODE SN 970